An environmental group estimates that the amount of pollution in a pond will decrease by 10% each month. To determine the concentration C, of pollutants (in parts per million) after (m) months, the group uses the function C(m)=23(0.95)^m
1) to the nearest tenth, what will be the concentration of pollutants in the pond in 5 years?
(2) What was the initial concentration of pollutants in the pond?

Answer by nerdybill(7384)

An environmental group estimates that the amount of pollution in a pond will decrease by 10% each month. To determine the concentration C, of pollutants (in parts per million) after (m) months, the group uses the function C(m)=23(0.95)^m
.
1) to the nearest tenth, what will be the concentration of pollutants in the pond in 5 years?
months in 5 years = 5*12 = 60 months
C(m)=23(0.95)^m
C(60)=23(0.95)^60
C(60)=23(.0461)
C(60)=1.06 parts per million
.
(2) What was the initial concentration of pollutants in the pond?
Initial concentration can be found by setting m=0
C(m)=23(0.95)^m
C(0)=23(0.95)^0
C(0)=23(1)
C(0)=23 parts per million
.